OpenStudy (anonymous):

Part 1 [3 points]: Find the solution to the equation 64(x – 3) = 42x. Part 2 [4 points]: Use complete sentences to explain the method used to solve this equation.

OpenStudy (campbell_st):

1. expand 64x - 192 = 42x 24x = 192 x = 8 2. to remove the brackets multiply each term in the brackets by 64 this gives 64x - 192 = 42x subtract 42x from both sides leaving 24x - 192 = 0 add 192 to both sides leaving 24x = 192 divide both sides by 24 giving the answer x = 8